3x+2y=22 2x+3y=23 using the simultaneous method find the values of x and y which satisfy the equation
Answer by ankor@dixie-net.com(12692)

3x+2y=22 2x+3y=23 using the simultaneous method find the values of x and y which satisfy the equation
:
3x + 2y = 22
2x + 3y = 23
:
Multiply eq one by 3 and eq two by 2 and you have:
9x + 6y = 66
4x + 6y = 46
-----------------Subtracting eliminates y
5x + 0y = 20
5x = 20
x = 20/5; divided both sides by 5
x = 4
:
Find y using equation one:
3x + 2y = 22
3(4) + 2y = 22
12 + 2y = 22
2y = 22 - 12, subtracted 12 from both sides
2y = 10
y = 10/2; divided both sides by 2
y = 5
:
:
Check solutions in equation two
2x + 3y = 23
2(4) + 3(5) = 23
8 + 15 = 23